# Cross-Modal Causal Intervention for Medical Report Generation

### v4

@@ IV-A4 Implementation Settings
We use the first three blocks of ResNet101 [50] to extract 1,024 feature maps, which are projected into 512 maps of size . The dimension of the transformer layers and the number of attention heads are fixed to 512 and 8, respectively. In the IU-Xray dataset (referred to as VLCI), both the encoder and decoder consist of 3 layers, whereas in the MIMIC-CXR dataset (referred to as VLCI), both the encoder and decoder comprise 6 layers, as illustrated in Table I. The variation in the number of layers can be attributed to the significantly larger size of the MIMIC-CXR dataset in comparison to the IU-Xray dataset. During pre-training, we utilize a dataset that combines IU-Xray and MIMIC-CXR datasets, resulting in 4,347 unique words. We perform fine-tuning on these two datasets separately with the same tokenizer as VLP. The batch size is set to 64 during pre-training and 16 during fine-tuning. In the pre-training stage, we adopt an image mask rate of 85% for MIM. The VLP is trained using the AdamW optimizer with a warm-up step of 10% of the total training steps, and the peak learning rate is set to 5e-4. The weight decay of the optimizer is set to 1e-2, and the total epochs are set to 30 in pre-training. In the fine-tuning stage, the model is fine-tuned using the Adam optimizer with an initial learning rate of 1e-5 and a weight decay of 5e-5 for 10 epochs on both the IU-Xray and MIMIC-CXR datasets.
@@ IV-B1 NLG Metric
As shown in Table III, our VLCI outperforms nearly all the MRG methods. Specifically, compared with the lightweight model MMTN [61], VLCI significantly improves the BLEU-1 metric by 1.9% on the IU-Xray dataset. Similarly, when compared with the heavyweight model RAMT [62], VLCI achieves a substantial 3.8% boost in the BLEU-1 metric on the MIMIC-CXR dataset. This significant improvement in the BLEU-1 metric highlights our method’s ability to select words more precisely. Notably, on the IU-Xray dataset, our tokenizer incorporates words from the MIMIC-CXR dataset, leading to even greater estimation challenges. To assess text fluency more effectively, we also consider the precise match of four consecutive words using the BLEU-4 metric. Our VLCI demonstrates a 1.5% improvement over MMTN on the IU-Xray dataset and a 0.6% improvement over RAMT on the MIMIC-CXR dataset. Notably, the longer descriptions in the MIMIC-CXR dataset present visual-linguistic data bias and significant spurious correlations among multiple words. Our VLCI effectively leverages cross-modal causal intervention to achieve performance improvement.
@@
Similarly, the METEOR metric considers synonyms and phrase reordering, emphasizing diverse matches of phrases and vocabulary. As indicated in Table III, VLCI achieves the best performance in terms of the METEOR metric on the IU-Xray dataset, but it falls short compared to MMTN on the MIMIC-CXR dataset. This suggests that there are some limitations in matching phrase and vocabulary diversity in our approach. By referring to Table I, we observe that both the best-performing MMTN and AlignTransformer benefit from knowledge supplementation and the assistance of a memory module, potentially enhancing the model’s ability to estimate synonyms more effectively. While our approach relies on learning semantics solely from the linear projection space of tokens, which may have limitations.
Moreover, CIDEr metrics emphasizes the importance of more significant and information-rich words in the generated text, rather than common words, posing a challenge for long sequence reports containing medical terminology. on the MIMIC-CXR dataset, VLCI outperforms two open-source non-knowledge-based methods (R2Gen and R2GenCMN) but falls behind knowledge-based methods (PPKED, MGSK, DCL). We speculate that domain-specific knowledge with specialized concepts can assist the model in generating more accurate disease descriptions and achieving superior performance, even if these descriptions may be less common. Our VLCI focuses on causal intervention within cross-modal data without relying on external knowledge, making it unable to accurately capture domain-specific terminology.
@@ IV-B2 CE Metric
The purpose of MRG is to alleviate the burden on medical professionals and provide precise diagnostic evidence. Therefore, the CE metric is more appropriate for evaluating the clinical significance of MRG methods compared to NLG metrics, as it specifically assesses the accuracy of abnormality classification. The CE metric is only applied to the MIMIC-CXR dataset because the label extractor (CheXpert) [69] is specially designed for MIMIC-CXR to obtain class labels. Compared with the lightweight R2Gen in Table III, VLCI demonstrates a remarkable improvement of 15.6% in Precision, 6.7% in Recall, and 12.5% in F1-Score. This validates that VLCI can provide a more accurate clinic diagnosis rather than merely generating fluent reports. Additionally, our VLCI outperforms the state-of-the-art method DCL in terms of Precision score and F1 score. Notably, when compared to several knowledge-based methods, our approach achieves a more efficient clinical evidence generation by relying solely on causal intervention without requiring additional knowledge assistance. Nevertheless, the Recall score is relatively lower compared to M2TR in Table III, indicating that the dataset may contain extreme categories. M2TR employs staged generation, which leverages the preliminary concepts generated in the first stage, enabling effective anomaly detection. However, this process can also lead to an increase in false positives, consequently reducing the Precision score.
@@ IV-C Qualitative Analysis
To further validate the assistance of causal intervention (i.e., alleviate the burden on medical professionals and provide precise diagnostic evidence) in our method, we extract 14 categories of labels from the reports generated by the baseline and VLCI, and evaluated their accuracy, as shown in Fig. 6. Our approach achieves significant performance improvements in all categories, particularly in “Edema” () and “Enlarge Cardiomediastinum” (). This is because our VLCI explores sufficient visual information and further produces more accurate and less biased descriptions by cross-modal causal intervention than the Transformer baseline. However, the estimation of some categories remains ambiguous, e.g., “Lung Opacity”. It reveals that VLCI can provide a comprehensive consideration of various radiologic signs to detect the abnormality but give less improvement for the single source abnormality. For example, whether “Edema” is caused by the heart has different radiologic signs, while the increase in lung density can be considered as “Lung Opacity”. Thus, VLCI can capture the abnormality with complex causes more effectively, where exists more spurious correlations. Besides, Fig. 6 shows the unavailability of causal intervention in independent abnormalities, e.g. “Support Devices”.
@@
We further conduct the qualitative analysis on the MIMIC-CXR dataset via three intuitive generated examples of the baseline and the VLCI in Fig. 7. In Fig. 7 (a), the reference report contains three abnormalities. However, the baseline model neglects ”pleural effusion” and ”consolidation,” while VLCI accurately identifies all abnormalities. This indicates that our VDM comprehensively captures all essential visual features, crucial for MRG. Fig. 7 (b) shows an example where the same visual region is simultaneously discovered by the baseline and the VLCI but leads to different descriptions. Our VLCI can accurately describe the heart, while the baseline is uncertain and even has a miscalculation of pneumonia. It shows that LDM can alleviate the semantic bias caused by word frequency in word embedding space. Fig. 7 (c) illustrates a complex causal graph, where “atelectasis” and “edema” could also be causes of “cardiomegaly.” However, the baseline fails to correctly consider the causes of “cardiomegaly” and erroneously captures these pieces of evidence. In contrast, VLCI leverages the causal intervention module to disentangle these confounders, enabling careful consideration of various pieces of evidence for an accurate judgment.
@@
In Fig. 7 (d), the ground truth indicates the presence of hydropneumothorax, a condition characterized by the simultaneous presence of gas and fluid in the chest, whereas pleural effusion contains only fluid. Although VLCI correctly identifies the presence of gas and fluid in the chest, leading to the diagnosis of pneumothorax, it erroneously estimates pleural effusion due to insufficient relevant knowledge. In this example, our method produces inaccurate text and fails to identify pneumonia and lung consolidation. Moreover, while VDM and LDM excel at recognizing visual and language concepts, detecting highly specialized concepts with latent relations not explicitly present in the data presents challenges.
@@ IV-D Ablation Studies
In Table IV, we adopt the same setting as ResNet101 to conduct the ablation experiments using DenseNet121 as the backbone. We perform VLP and fine-tuning with causal intervention on IU-Xray and MIMIC-CXR datasets. Due to the small scale of the IU-Xray dataset and the simplicity of report content, the persuasiveness of the CIDEr metric on this dataset is limited. We evaluated BLEU-4, Rouge-L, and METEOR on the IU-Xray dataset, and the results indicate that our method still achieves a significant improvement with DenseNet121. However, we observed that the performance of models using DenseNet121 as the backbone is notably inferior to those using ResNet101. We speculate that while DenseNet121 is efficient in feature extraction due to its dense connections, it may not always provide the best performance for tasks involving specific types of medical images or requiring deeper feature extraction capabilities. Additionally, we conducted experiments on the larger MIMIC-CXR dataset, using BLEU-4, CIDEr, and METEOR as evaluation metrics. The experimental results further confirm our speculations and validate the effectiveness of our approach.
| Method on IU-Xray | BLEU-4 | Rouge-L | METEOR |
| ResNet101 | 0.148 | 0.345 | 0.180 |
| ResNet101 + VLCI | 0.190 (+0.042) | 0.394 (+0.049) | 0.210 (+0.03) |
| DenseNet121 | 0.164 | 0.361 | 0.183 |
| DenseNet121 + VLCI | 0.176 (+0.012) | 0.394 (+0.033) | 0.195 (+0.012) |
| Method on MIMIC-CXR | BLEU-4 | CIDEr | METEOR |
| ResNet101 | 0.101 | 0.130 | 0.135 |
| ResNet101 + VLCI | 0.119 (+0.018) | 0.190 (+0.06) | 0.150 (+0.015) |
| DenseNet121 | 0.091 | 0.111 | 0.127 |
| DenseNet121 + VLCI | 0.113 (+0.022) | 0.134 (+0.023) | 0.149 (+0.022) |
IV-D2 Effectiveness of VLP
In Table V, we conduct ablation experiments to assess the impact of masking ratios on model performance, and the results are presented. Our VLP model achieved the best performance with a higher masking ratio of 85%, which is in contrast to the optimal masking ratio of 75% reported by MAE [48]. We attribute this difference to the cross-modal information correlations, where the masked information can be reconstructed by visible features from both language and images. Furthermore, VLP tends to learn general features from the masked modality at higher masking ratios, while distinguishable features can be extracted by the complete information from another modality. To explore whether increasing the masking ratio further would further improve the performance, we experimented with a higher masking ratio of 95%. However, the decreased results in Table V indicate that this approach leads to excessive information loss.
| Masking Ratio | BLEU-1 | BLEU-4 | CIDEr | Rouge-L |
| Baseline | 0.433 | 0.148 | 0.501 | 0.345 |
| w/ 75% | 0.450 | 0.160 | 0.486 | 0.360 |
| w/ 85% | 0.452 | 0.161 | 0.522 | 0.351 |
| w/ 95% | 0.432 | 0.153 | 0.460 | 0.346 |
@@ IV-D2 Effectiveness of VLP
| Baseline | 0.148 | 0.501 | 0.345 |
| w/ MAE | 0.154 | 0.486 | 0.360 |
| w/ VisualGPT | 0.159 | 0.497 | 0.374 |
| w/ VLP (MIM) | 0.162 | 0.602 | 0.362 |
| w/ VLP (PLM) | 0.165 | 0.538 | 0.365 |
| w/ VLP (PLM+MIM) | 0.160 | 0.431 | 0.364 |
| w/ VLP (PLM)* | 0.151 | 0.399 | 0.349 |
| w/ VLP (PLM+MIM)* | 0.161 | 0.522 | 0.351 |
| w/ VLP (PLM+MIM+Degradation)* (Ours) | 0.170 | 0.631 | 0.363 |
In Table VI, we make a comparison with different pre-training methods. It shows that the cross-modal pre-training method has a more robust representation ability than the MIM with single-modality. However, the Rouge-L metric in VisualGPT surpasses ours, possibly due to its exclusive pre-training of the text decoder, enabling more concentrated learning of the intricate structure in medical reports. Additionally, our cross-modal pre-training achieves comparable performance to the PLM model that only fine-tunes the encoder, while ours fine-tunes the whole model with fewer epochs. Moreover, our VLP adopts the degradation images as input, which facilitates the extraction of visual details in the downstream task.
@@
| Baseline | 0.148 | 0.501 | 0.345 | |
| Baseline (w/ VDM) | 0.160 | 0.521 | 0.364 | |
| Baseline (w/ LDM) | 0.155 | 0.509 | 0.361 | |
| Baseline (w/ VDM&LDM) | 0.163 | 0.544 | 0.361 | |
| R2Gen | 0.165 | 0.493 | 0.360 | |
| R2Gen (w/ VDM) | 0.171 | 0.553 | 0.370 | |
| IU-Xray | R2Gen (w/ LDM) | 0.166 | 0.546 | 0.360 |
| R2Gen (w/ VDM&LDM) | 0.173 | 0.628 | 0.368 | |
| Baseline (w/ VLP) | 0.170 | 0.631 | 0.363 | |
| Baseline (w/ VLP&VDM) | 0.174 | 0.523 | 0.374 | |
| Baseline (w/ VLP&LDM) | 0.178 | 0.573 | 0.378 | |
| VLCI | 0.190 | 0.592 | 0.394 | |
| Baseline | 0.101 | 0.130 | 0.270 | |
| Baseline (w/ VDM) | 0.103 | 0.144 | 0.272 | |
| Baseline (w/ LDM) | 0.069 | 0.071 | 0.224 | |
| Baseline (w/ VDM&LDM) | 0.070 | 0.074 | 0.230 | |
| R2Gen | 0.103 | 0.168 | 0.278 | |
| R2Gen (w/ VDM) | 0.106 | 0.171 | 0.277 | |
| MIMIC-CXR | R2Gen (w/ LDM) | 0.091 | 0.136 | 0.256 |
| R2Gen (w/ VDM&LDM) | 0.100 | 0.143 | 0.264 | |
| Baseline (w/ VLP) | 0.106 | 0.151 | 0.278 | |
| Baseline (w/ VLP&VDM) | 0.110 | 0.177 | 0.280 | |
| Baseline (w/ VLP&LDM) | 0.115 | 0.157 | 0.277 | |
| VLCI | 0.119 | 0.190 | 0.280 |
Furthermore, in Table VII, Baseline is significantly worse than baseline on the MIMIC-CXR dataset, e.g., 0.101 0.070 for BLEU-4, while still keeping performance improvement on IU-Xray dataset. This validates the significant feature complexity from the large-scale MIMIC-CXR dataset leads to unstable probability distribution estimation with causal intervention. Meanwhile, we find that the VLP can substantially boost the performance of the baseline, e.g., 0.148 0.170, 0.101 0.106 for BLEU-4 on IU-Xray and MIMIC-CXR datasets, respectively. The improvement is caused by the learned comprehensive concepts and context in the pre-training and the cross-modal features alignment stage, which shows the importance of VLP. Similarly, The Rough-L is also barely improved due to the features’ complexity and long sequence from the MIMIC-CXR dataset. For example, although AlignTransformer achieves the same score of the Rough-L as CA on the MIMIC-CXR, it outperforms CA on all other metrics.

### v5

@@ IV-A4 Implementation Settings
We use the first three blocks of ResNet101 [67] to extract 1,024 feature maps, which are projected into 512 maps of size . The dimension of the transformer layers and the number of attention heads are fixed to 512 and 8, respectively. In the IU-Xray dataset (referred to as CMCRL3), both the encoder and decoder consist of 3 layers, whereas in the MIMIC-CXR dataset (referred to as CMCRL6), both the encoder and decoder comprise 6 layers, as illustrated in Table I. The variation in the number of layers can be attributed to the significantly larger size of the MIMIC-CXR dataset in comparison to the IU-Xray dataset. During pre-training, we utilize a dataset that combines IU-Xray and MIMIC-CXR datasets, resulting in 4,347 unique words. We perform fine-tuning on these two datasets separately with the same tokenizer as RadCARE. The batch size is set to 64 during pre-training and 16 during fine-tuning. In the pre-training stage, we adopt an image mask rate of 85% for masked image restoration. The RadCARE is trained using the AdamW optimizer with a warm-up step of 10% of the total training steps, and the peak learning rate is set to 5e-4. The weight decay of the optimizer is set to 1e-2, and the total epochs are set to 30 in pre-training. In the fine-tuning stage, the model is fine-tuned using the Adam optimizer with an initial learning rate of 1e-5 and a weight decay of 5e-5 for 10 epochs on both the IU-Xray and MIMIC-CXR datasets.
@@ IV-B1 NLG Metric
As shown in Table III, our CMCRL outperforms nearly all the RRG methods. Specifically, compared with the lightweight model MMTN [79], CMCRL3 significantly improves the BLEU-1 metric by 1.9% on the IU-Xray dataset. Similarly, when compared with the heavyweight model RAMT [80], CMCRL6 achieves a substantial 3.8% boost in the BLEU-1 metric on the MIMIC-CXR dataset. This significant improvement in the BLEU-1 metric highlights our method’s ability to select words more precisely. Notably, on the IU-Xray dataset, our tokenizer incorporates words from the MIMIC-CXR dataset, leading to even greater estimation challenges. To assess text fluency more effectively, we also consider the precise match of four consecutive words using the BLEU-4 metric. Our CMCRL demonstrates a 1.5% improvement over MMTN on the IU-Xray dataset and a 0.6% improvement over RAMT on the MIMIC-CXR dataset. Notably, the longer descriptions in the MIMIC-CXR dataset present visual-linguistic data bias and significant spurious correlations among multiple words. Our CMCRL effectively leverages cross-modal causal intervention to achieve performance improvement.
@@
Similarly, the METEOR metric considers synonyms and phrase reordering, emphasizing diverse matches of phrases and vocabulary. As indicated in Table III, CMCRL achieves the best performance in terms of the METEOR metric on the IU-Xray dataset, but it falls short compared to MMTN on the MIMIC-CXR dataset. This suggests that there are some limitations in matching phrase and vocabulary diversity in our approach. By referring to Table I, we observe that both the best-performing MMTN and AlignTransformer benefit from knowledge supplementation and the assistance of a memory module, potentially enhancing the model’s ability to estimate synonyms more effectively. While our approach relies on learning semantics solely from the linear projection space of tokens, which may have limitations.
@@ IV-B2 CE Metric
The purpose of RRG is to alleviate the burden on radiological professionals and provide precise diagnostic evidence. Therefore, the CE metric is more appropriate for evaluating the clinical significance of RRG methods compared to NLG metrics, as it specifically assesses the accuracy of abnormality classification. The CE metric is only applied to the MIMIC-CXR dataset because the label extractor (CheXpert) [87] is specially designed for MIMIC-CXR to obtain class labels. Compared with the lightweight R2Gen in Table III, CMCRL demonstrates a remarkable improvement of 15.6% in Precision, 6.7% in Recall, and 12.5% in F1-Score. This validates that CMCRL can provide a more accurate clinic diagnosis rather than merely generating fluent reports. Additionally, our CMCRL outperforms the state-of-the-art method DCL in terms of Precision score and F1 score. Notably, when compared to several knowledge-based methods, our approach achieves a more efficient clinical evidence generation by relying solely on causal intervention without requiring additional knowledge assistance. Nevertheless, the Recall score is relatively lower compared to M2TR in Table III, indicating that the dataset may contain extreme categories. M2TR employs staged generation, which leverages the preliminary concepts generated in the first stage, enabling effective anomaly detection. However, this process can also lead to an increase in false positives, consequently reducing the Precision score.
@@ IV-C1 Abnormalities Detection
To further validate the assistance of causal intervention (i.e., alleviate the burden on radiological professionals and provide precise diagnostic evidence) in our method, we extract 14 categories of labels from the reports generated by the baseline and CMCRL, and evaluated their accuracy, as shown in Fig. 8. Our approach achieves significant performance improvements in all categories, particularly in “Edema” () and “Enlarge Cardiomediastinum” (). This is because our CMCRL explores sufficient visual information and further produces more accurate and less biased descriptions by cross-modal causal intervention than the Transformer baseline. However, the estimation of some categories remains ambiguous, e.g., “Lung Opacity”. It reveals that CMCRL can provide a comprehensive consideration of various radiologic signs to detect the abnormality but give less improvement for the single source abnormality. For example, whether “Edema” is caused by the heart has different radiologic signs, while the increase in lung density can be considered as “Lung Opacity”. Thus, CMCRL can capture the abnormality with complex causes more effectively, where exists more spurious correlations. Besides, Fig. 8 shows the unavailability of causal intervention in independent abnormalities, e.g. “Support Devices”.
IV-C2 Causal Consistency
To assess whether CMCRL utilized appropriate visual evidence for its inference, we develop a questionnaire and distributed it to human experts for evaluation. We select 10 data categories, each containing 10 samples, resulting in a total of 100 instances. Visual attention maps are generated based on the descriptions in the reports, where attention is captured through the cumulative response of entire sentences. These visualizations are compiled into a questionnaire (as shown in Figure 9 (a-c)) and distributed to 12 experienced radiological professionals. They evaluate the plausibility of the results based on the original images, attention maps, and corresponding sentences.
As shown in Fig 9 (d), our approach achieves an overall reasonableness rating of 76.3%, indicating that most of the samples are accepted by experts. Moreover, due to the low image resolution (224*224), the experts find it challenging to make judgments. Furthermore, the results for Cardiomegaly, Pneumothorax, and Support Devices exhibited significant consistency with human priors. This underscores the effectiveness of our approach in uncovering latent causal relations rather than relying on spurious correlations for judgment. Additionally, we intend to further advance the field in interpretable and trustworthy reasoning, with the goal of leveraging explicit causal relations for inference in the future.
IV-C3 Case Study
@@
We further conduct the qualitative analysis on the MIMIC-CXR dataset via three intuitive generated examples of the baseline and the CMCRL in Fig. 10. In Fig. 10 (a), the reference report contains four abnormalities. However, the baseline model neglects all abnormalities, while CMCRL accurately identifies all abnormalities. This indicates that our VDM comprehensively captures all essential visual features, crucial for RRG. Fig. 10 (b) shows an example where the same visual region is simultaneously discovered by the baseline and the CMCRL but leads to different descriptions. Our CMCRL can accurately describe the heart, while the baseline is uncertain and even has a miscalculation of pneumonia. It shows that LDM can alleviate the semantic bias caused by word frequency in word embedding space. Fig. 10 (c) illustrates a complex causal graph, where “atelectasis” and “edema” could also be causes of “cardiomegaly.” However, the baseline fails to correctly consider the causes of “cardiomegaly” and erroneously captures these pieces of evidence. In contrast, CMCRL leverages the causal intervention module to disentangle these confounders, enabling careful consideration of various pieces of evidence for an accurate judgment.
@@
In Fig. 10 (d), the ground truth indicates the presence of hydropneumothorax, a condition characterized by the simultaneous presence of gas and fluid in the chest, whereas pleural effusion contains only fluid. Although CMCRL correctly identifies the presence of gas and fluid in the chest, leading to the diagnosis of pneumothorax, it erroneously estimates pleural effusion due to insufficient relevant knowledge. In this example, our method produces inaccurate text and fails to identify pneumonia and lung consolidation. Moreover, while VDM and LDM excel at recognizing visual and language concepts, detecting highly specialized concepts with latent relations not explicitly present in the data presents challenges.
Method on IU-Xray | BLEU-4 | Rouge-L | METEOR |
ResNet101 | 14.8 | 34.5 | 18.0 |
ResNet101 + VLCI | 19.0 (+4.2) | 39.4 (+4.9) | 21.0 (+3.0) |
DenseNet121 | 16.4 | 36.1 | 18.3 |
DenseNet121 + VLCI | 17.6 (+1.2) | 39.4 (+3.3) | 19.5 (+1.2) |
Method on MIMIC-CXR | BLEU-4 | CIDEr | METEOR |
ResNet101 | 10.1 | 13.0 | 13.5 |
ResNet101 + VLCI | 11.9 (+1.8) | 19.0 (+6) | 15.0 (+1.5) |
DenseNet121 | 9.1 | 11.1 | 12.7 |
DenseNet121 + VLCI | 11.3 (+2.2) | 13.4 (+2.3) | 14.9 (+2.2) |
Masking Ratio | BLEU-1 | BLEU-4 | CIDEr | Rouge-L |
Baseline | 43.3 | 14.8 | 50.1 | 34.5 |
w/ 75% | 45.0 | 16.0 | 48.6 | 36.0 |
w/ 85% | 45.2 | 16.1 | 52.2 | 35.1 |
w/ 95% | 43.2 | 15.3 | 46.0 | 34.6 |
@@ IV-D1 Effectiveness of VLCI
In Table IV, we adopt the same setting as ResNet101 to conduct the ablation experiments using DenseNet121 as the backbone. We perform RadCARE and fine-tuning with causal intervention on IU-Xray and MIMIC-CXR datasets. Due to the small scale of the IU-Xray dataset and the simplicity of report content, the persuasiveness of the CIDEr metric on this dataset is limited. We evaluated BLEU-4, Rouge-L, and METEOR on the IU-Xray dataset, and the results indicate that our method still achieves a significant improvement with DenseNet121. However, we observed that the performance of models using DenseNet121 as the backbone is notably inferior to those using ResNet101. We speculate that while DenseNet121 is efficient in feature extraction due to its dense connections, it may not always provide the best performance for tasks involving specific types of radiological images or requiring deeper feature extraction capabilities. Additionally, we conducted experiments on the larger MIMIC-CXR dataset, using BLEU-4, CIDEr, and METEOR as evaluation metrics. The results further confirm our speculations and validate the effectiveness of our approach.
@@
Baseline | 14.8 | 50.1 | 34.5 |
w/ MAE | 15.4 | 48.6 | 36.0 |
w/ VisualGPT | 15.9 | 49.7 | 37.4 |
w/ MIM | 16.2 | 60.2 | 36.2 |
w/ TEXT | 16.5 | 53.8 | 36.5 |
w/ TEXT+IMAGE | 16.0 | 43.1 | 36.4 |
w/ TEXT* | 15.1 | 39.9 | 34.9 |
w/ TEXT+IMAGE* | 16.1 | 52.2 | 35.1 |
w/ RadCARE* (Ours) | 17.0 | 63.1 | 36.3 |
@@
Baseline | 14.8 | 50.1 | 34.5 | |
Baselinew◆ (w/ VDM) | 16.0 | 52.1 | 36.4 | |
Baselinew∙ (w/ LDM) | 15.5 | 50.9 | 36.1 | |
Baselinew◆∙ (w/ VDM&LDM) | 16.3 | 54.4 | 36.1 | |
R2Gen | 16.5 | 49.3 | 36.0 | |
R2Genw◆ (w/ VDM) | 17.1 | 55.3 | 37.0 | |
IU-Xray | R2Genw∙ (w/ LDM) | 16.6 | 54.6 | 36.0 |
R2Genw◆∙ (w/ VDM&LDM) | 17.3 | 62.8 | 36.8 | |
Baselinew★ (w/ RadCARE) | 17.0 | 63.1 | 36.3 | |
Baselinew★◆ (w/ RadCARE&VDM) | 17.4 | 52.3 | 37.4 | |
Baselinew★∙ (w/ RadCARE&LDM) | 17.8 | 57.3 | 37.8 | |
CMCRL | 19.0 | 59.2 | 39.4 | |
Baseline | 10.1 | 13.0 | 27.0 | |
Baselinew◆ (w/ VDM) | 10.3 | 14.4 | 27.2 | |
Baselinew∙ (w/ LDM) | 6.9 | 7.1 | 22.4 | |
Baselinew◆∙ (w/ VDM&LDM) | 7.0 | 7.4 | 23.0 | |
R2Gen | 10.3 | 16.8 | 27.8 | |
R2Genw◆ (w/ VDM) | 10.6 | 17.1 | 27.7 | |
MIMIC-CXR | R2Genw∙ (w/ LDM) | 9.1 | 13.6 | 25.6 |
R2Genw◆∙ (w/ VDM&LDM) | 10.0 | 14.3 | 26.4 | |
Baselinew★ (w/ RadCARE) | 10.6 | 15.1 | 27.8 | |
Baselinew★◆ (w/ RadCARE&VDM) | 11.0 | 17.7 | 28.0 | |
Baselinew★∙ (w/ RadCARE&LDM) | 11.5 | 15.7 | 27.7 | |
CMCRL | 11.9 | 19.0 | 28.0 |
IV-D2 Effectiveness of RadCARE
In Table V, we conduct ablation experiments to assess the impact of masking ratios on model performance, and the results are presented. Our RadCARE model achieved the best performance with a higher masking ratio of 85%, which is in contrast to the optimal masking ratio of 75% reported by MAE [17]. We attribute this difference to the cross-modal information correlations, where the masked information can be reconstructed by visible features from both language and images. Furthermore, RadCARE tends to learn general features from the masked modality at higher masking ratios, while distinguishable features can be extracted by the complete information from another modality. To explore whether increasing the masking ratio further would further improve the performance, we experimented with a higher masking ratio of 95%. However, the decreased results in Table V indicate that this approach leads to excessive information loss.
In Table VI, we make a comparison with different pre-training methods. It shows that the cross-modal pre-training method has a more robust representation ability than the masked image restoration with single-modality. However, the Rouge-L metric in VisualGPT surpasses ours, possibly due to its exclusive pre-training of the text decoder, enabling more concentrated learning of the intricate structure in radiological reports. Additionally, our cross-modal pre-training via postfix text generation task achieves comparable performance to the model that only fine-tunes the encoder, while ours fine-tunes the whole model with fewer epochs. Moreover, our RadCARE adopts the degradation images as input, which facilitates the extraction of visual details in the downstream task.
Furthermore, in Table VII, Baselinew◆∙ is significantly worse than baseline on MIMIC-CXR, e.g., 10.1 7.0 for BLEU-4, while still keeping performance improvement on IU-Xray dataset. This validates the significant feature complexity from the large-scale MIMIC-CXR dataset leads to unstable probability distribution estimation with causal intervention. Meanwhile, the RadCARE can substantially boost the performance of the baseline, e.g., 14.8 17.0, 10.1 10.6 for BLEU-4 on IU-Xray and MIMIC-CXR datasets, respectively. The improvement is caused by the learned comprehensive concepts and context in the pre-training and the cross-modal features alignment stage, which shows the importance of RadCARE. Similarly, The Rough-L is also barely improved due to the features’ complexity and long sequence from the MIMIC-CXR dataset. For example, although AlignTransformer achieves the same score of the Rough-L as CA on the MIMIC-CXR, it outperforms CA on all other metrics.